OpenStudy (anonymous):

MEDAL WILL BE REWARDED: the midpoint of CD is (6,5). if the cordinates of point C are (3,8), find the cordinates of point D.....i know the midpoint and distance formula but not sure how to answer this. The answer is (9,2).

jimthompson5910 (jim_thompson5910):

You need to solve these two equations \[\Large \frac{3+x}{2} = 6\] \[\Large \frac{8+y}{2} = 5\] to find the point D (x,y)

OpenStudy (anonymous):

Thank you so much :)

jimthompson5910 (jim_thompson5910):

you're welcome